# Concurrent Composition for Differentially Private Continual Mechanisms

### v3

@@ 3. Continual Mechanisms
Let $\mathcal{M}$ be a CM and $\mathcal{I}$ be an identifier for $\mathcal{M}$. A verifier $\mathcal{V}$ with a verification function $f$, denoted by $\mathcal{V}[f]$, stands between an adversary $\mathcal{A}$ and the interactive mechanism $\mathcal{I}\!\circ^{*}\!\mathcal{M}$. Upon receiving a message from $\mathcal{A}$, $\mathcal{V}[f]$ applies $f$ to verify whether the sequence of pairs sent by the adversary up to that point satisfies a set of properties. Comparing the sequences formed by the first and second elements of the pairs, $f$ can be used to encode a neighboring property. If the pair sequence satisfies the properties, $\mathcal{V}[f]$ forwards $\mathcal{A}$’s message to $\mathcal{I}\!\circ^{*}\!\mathcal{M}$; otherwise, it halts the communication. The formal definitions of identifiers, verification functions, and verifiers are given below.
@@
Since the set of all possible messages does not exist, we assume that all message spaces considered in this paper are subsets of some universal set $\mathcal{U}$. For instance, in [11], the universal set is taken to be $\{0,1\}^{*}$. The domain of verification functions is $\mathcal{U}^{*}$.
@@
For a verification function $f$, a verifier $\mathcal{V}[f]$ is an IPM with the initial state space $S_{\mathcal{V}[f]}^{\text{init}}=\{()\}$, where $()$ represents an empty sequence. Since the initial state space of $\mathcal{V}[f]$ is a singleton, we denote an instance of $\mathcal{V}[f]$ with the initial state $()$ as $\mathcal{V}[f]$. $\mathcal{V}[f]$ stands between two IMs, referred to as the left and right mechanisms. The state of $\mathcal{V}[f]$ consists of the sequence of messages it has received from the left mechanism so far. Upon receiving a message $m$ from the left mechanism, $\mathcal{V}[f]$ appends $m$ to its state and applies the verification function $f$. If the sequence is $f$-valid, then $\mathcal{V}[f]$ forwards the message $m$ to the right mechanism; otherwise, it halts the communication. $\mathcal{V}[f]$ forwards any message from the right mechanism unchanged to the left mechanism.
@@
Let $S_{\mathcal{M}}^{\text{init}}=\{s^{\text{init}}\}$ be the initial state space of the continual mechanism $\mathcal{M}$ in Definition 3.7. By definition, the initial state spaces of $\mathcal{I}$ and $\mathcal{V}[f]$ are $\{0,1\}$ and $\{()\}$, respectively. Thus, the initial state space of $\mathcal{V}[f]\!\circ^{*}\!\mathcal{I}\!\circ^{*}\!\mathcal{M}$ is $\left\{((),0,s^{\text{init}}),((),1,s^{\text{init}})\right\}$. Define $\sim_{01}$ as a neighboring on this set, where only $((),0,s^{\text{init}})\sim_{01}((),1,s^{\text{init}})$. Comparing Definition 3.7 and Definition 2.13 leads to the following corollary:
@@ 4. Concurrent Composition of Continual Mechanisms
Concurrent Composition of IMs. Vadhan and Wang [23] extended the definition of the composition of NIMs to the concurrent composition of IMs, where an adversary interacts with $k$ interactive mechanisms $\mathcal{M}_{1},\dots,\mathcal{M}_{k}$ at the same time. In concurrent composition, the adversary generates a message and determines the receiver of this message (one of the $\mathcal{M}_{i}$s) adaptively based on all the previous messages exchanged between the adversary and the mechanisms $M_{1},\dots,M_{k}$. To formalize this, Vadhan and Wang [23] introduced an interactive mechanism ${\textrm{ConComp}}[\mathcal{M}_{1},\dots,\mathcal{M}_{k}]$, which internally runs $\mathcal{M}_{1},\dots,\mathcal{M}_{k}$ and, upon receiving a message $(q,j)$, forwards the query $q$ to $\mathcal{M}_{j}$ and returns its response.
@@
Vadhan and Wang [23] show that if each IM $\mathcal{M}_{i}$ is $\epsilon_{i}$-DP w.r.t $\sim_{i}$, and if the composition of NIMs $\mathrm{RR}_{\epsilon_{1},0},\dots,\mathrm{RR}_{\epsilon_{k},0}$ is $(\epsilon,\delta)$-DP, then ${\textrm{ConComp}}[\mathcal{M}_{1},\dots,\mathcal{M}_{k}]$ is also $(\epsilon,\delta)$-DP against an adaptive adversary. Lyu [19] extends this result for $(\epsilon_{i},\delta_{i})$-DP IMs.
In the concurrent composition of IMs, queries asked to each mechanism are chosen adaptively, not only based on that mechanism’s previous responses but also using the responses of the other mechanisms up to that point. The results of [23] and [19] show that this concurrency and additional adaptivity do not compromise privacy guarantees. As a result, all $(\epsilon,\delta)$-DP composition theorems for NIMs, such as the basic [6], advanced [9], and optimal [16] composition theorems, can also be applied to the concurrent composition of IMs.
@@
To extend Lyu’s result on the privacy guarantee of the concurrent composition of IMs to the concurrent composition of CMs, we want to use the following intermediate lemma from their paper: For every $(\epsilon,\delta)$-DP interactive mechanism $\mathcal{M}$ and every pair of neighboring initial states $s$ and $s^{\prime}$ for $\mathcal{M}$, there exists an IPM $\mathcal{P}$ such that for each $b\in\{0,1\}$, the IMs $\mathcal{P}\!\circ^{*}\!\mathrm{RR}_{\epsilon,\delta}(b)$ and $\mathcal{M}(b)$ are equivalent. However, as we discuss in Section 4.2, there are hidden assumptions in the proof of this lemma. Additionally, we show there that the construction of $\mathcal{P}$ in [23], for $\epsilon_{i}$-DP IMs, satisfies a property, required for a later proof.
@@
We extend the definition of ConComp in [23] and introduce ExtConComp that allows the creation of new mechanisms in addition to queries to existing ones. A creation query $(\mathcal{M},s,\epsilon,\delta,f)$ is a message that requests the creation of a continual mechanism $\mathcal{M}$ with the initial state $s$ and asserts that $\mathcal{M}$ is $(\epsilon,\delta)$-DP w.r.t. the verification function $f$. The mechanism ExtConComp accepts creation queries and messages of the form $(q,i)$, where $q$ is a query for the $i$-th created mechanism. The formal definition of ExtConComp is as follows:
@@
Next, we formalize these shared constraints by defining suitable message sequences. Later, when defining specific verification functions $f$ for different concurrent compositions, we set $f(m_{1},\dots,m_{k})=\bot$ whenever the sequence of messages $m_{1},\dots,m_{k}$ is not suitable.
@@ 4.2. $(\epsilon,\delta)$-DP IMs and $\mathrm{RR}_{\epsilon,\delta}$
Lyu [19] shows that for every $(\epsilon,\delta)$-DP IM $\mathcal{M}$ and every pair of neighboring initial states $s_{0}$ and $s_{1}$, there exists an IPM $\mathcal{T}$ (with a single initial state) such that for each $b\in\{0,1\}$, the mechanisms $\mathcal{T}\!\circ^{*}\!\mathrm{RR}_{\epsilon,\delta}(b)$ and $\mathcal{M}(s_{b})$ are identical. Although not explicitly stated, the construction of $\mathcal{T}$ proposed in [19] requires the answer distributions of $\mathcal{M}(s_{0})$ and $\mathcal{M}(s_{1})$ to be discrete. This assumption is formally defined as:
@@
An IM $\mathcal{M}$ has discrete answer distributions if $A_{\mathcal{M}}$ is countable, and for every initial state $s$, integer $t\in\mathbb{N}$, history $(q_{1},a_{1},\dots,q_{t-1},a_{t-1})\in(Q_{\mathcal{M}}\times A_{\mathcal{M}})^{t-1}$, and new query $q_{t}\in Q_{\mathcal{M}}$, the distribution of $\mathcal{M}(s)$’s answer in respond to $q_{t}$ conditioned on having the history $(q_{1},\dots,a_{t-1})$ has a probability mass function (PMF).
@@
Lyu [19] additionally assumes that there exists a pre-fixed upper bound $c$ on the number of queries $\mathcal{M}$ answers. This assumption stems from their use of a backward-defined control function. By employing the limit control functions which we introduce in Section 8, one can modify their proof to remove this additional assumption. Moreover, using the formalization of Section 8 to define recursive functions, we can relax the requirement that the answer set is finite to allow countably infinite answer sets.
@@
Lemma 4.6 ([19]).
@@ 4.3. Concurrent Composition of CMs $\mathcal{M}_{1},\dots,\mathcal{M}_{k}$
In the concurrent composition of IMs, queries to each mechanism are chosen adaptively, not only based on that mechanism’s previous responses but also using the responses of the other mechanisms up to that point. By definition, the choice of neighboring initial states for the interactive mechanisms is non-adaptive. However, in the concurrent composition of CMs, the choice of differences between query sequences, i.e., the neighboring query sequences is adaptive, based on not only the responses of the corresponding mechanism but also the responses of other mechanisms. (Note that, unlike IMs, CMs inherently have a single fixed initial state; hence, no choice of initial state is involved.) Despite this additional adaptivity, in the following theorem, we show that the privacy results of [19] for ${\textrm{ConComp}}[\mathcal{M}_{1},\dots,\mathcal{M}_{k}]$ also hold for $\mathcal{V}[f^{\alpha_{1},\dots,\alpha_{k}}]\!\circ^{*}\!\mathcal{I}\!\circ^{*}\!{\textrm{ExtConComp}}$.
@@ 4.4. Concurrent Composition of CMs with Fixed Parameters
We refer to the $f^{\epsilon_{1},\delta_{1},\dots,\epsilon_{k},\delta_{k}}$-concurrent composition of CMs as the concurrent composition of CMs with privacy parameters $\left((\epsilon_{i},\delta_{i})\right)_{i=1}^{k}$. A slight adaptation of the proof of Theorem 4.9 implies the following theorem. For convenience, in the rest of the paper, we represent a pair of identical creation queries $(\mathcal{M},s,\epsilon,\delta,f)$ by $(\mathcal{M},s,\epsilon,\delta,f)^{2}$.
@@ 5.2. Composition Theorem
In the proofs of the previous composition theorems (namely, Theorems 4.9 and 4.11), each created CM $\mathcal{M}_{j}^{\prime}$ with privacy parameters $(\epsilon_{j}^{\prime},\delta_{j}^{\prime})$ is simulated by the IPM $\mathcal{T}_{j}$ from Lemma 4.6. For each $b\in\{0,1\}$, the mechanism $\mathcal{T}_{j}$ interacts with $\mathrm{RR}_{\epsilon_{j}^{\prime},\delta_{j}^{\prime}}(b)$ once and simulates $\mathcal{V}[f_{j}^{\prime}]\!\circ^{*}\!\mathcal{I}(b)\!\circ^{*}\!\mathcal{M}_{j}^{\prime}$ identically. Next, we present a new construction for $\mathcal{T}_{j}$, where it interacts with $\mathrm{IRR}_{\epsilon_{j}^{\prime},\delta_{j}^{\prime}}$ instead of $\mathrm{RR}_{\epsilon_{j}^{\prime},\delta_{j}^{\prime}}$. This construction satisfies additional properties needed in the proof of Theorem 5.4. In Section 8, we prove the following lemma, which directly implies the desired IPM as stated in Corollary 5.7.
@@
Since the parameters $\delta_{j}^{\prime}$ are chosen adaptively, the composition of $\mathrm{RR}_{0,\delta_{1}^{\prime}},\mathrm{RR}_{0,\delta_{2}^{\prime}},\dots$ is an instance of the filter composition of NIMs [21]. We do not state this composition in terms of filters and define it using a verification function $f^{\delta^{\prime}}_{\mathrm{RR}}$ that only allows the creation of randomized response mechanisms $\mathrm{RR}_{0,\delta_{j}^{\prime}}$ and returns $\bot$ if $1-\prod(1-\delta_{j}^{\prime})>\delta^{\prime}$.
@@
The following lemma is proved in Section 9.  [^6] By setting $\epsilon_{1}=\epsilon_{2}=\dots=0$ in Theorem 2 of [25], it follows that for every $\epsilon>0$, $\delta^{\prime}>0$, and $0\leq\delta^{\prime\prime}\leq 1$, the composition of randomized response mechanisms $\mathrm{RR}_{0,\delta_{1}},\mathrm{RR}_{0,\delta_{2}},\dots$ with adaptively chosen privacy parameters $\delta_{1},\delta_{2},\dots$ satisfying $\sum_{i}\delta_{i}\leq\delta^{\prime\prime}$ guarantees $(\epsilon,\delta^{\prime}+\delta^{\prime\prime})$-DP. Taking $\epsilon$ and $\delta^{\prime}$ to zero, we can conclude that this composition is $(0,\delta^{\prime\prime})$-DP. To get tighter bounds, we relax the requirement $\sum_{i}\delta_{i}\leq\delta^{\prime\prime}$ to $1-\prod_{i}(1-\delta_{i})\leq\delta^{\prime\prime}$ and prove the composition is still $(0,\delta^{\prime\prime})$-DP.
@@
For every $\epsilon\geq 0$ and $0\leq\delta,\delta^{\prime}\leq 1$, if ${\textrm{Comp}}(\mathrm{RR}_{\epsilon_{1},0},\dots,\mathrm{RR}_{\epsilon_{k},0})$ is $(\epsilon,\delta)$-DP, then the $f^{\epsilon_{1},\dots,\epsilon_{k}}_{\infty,\delta^{\prime},RR}$-concurrent composition of CMs is $(\epsilon,\delta+\delta^{\prime})$-DP.
@@
It remains to prove that the property (2) holds. The assumption that ${\textrm{Comp}}(\mathrm{RR}_{\epsilon_{1},0},\dots,\mathrm{RR}_{\epsilon_{k},0})$ is $(\epsilon,\delta)$-DP is equivalent to that ExtConComp is $(\epsilon,\delta)$-DP w.r.t. the verification function $f^{\beta_{1},\dots,\beta_{k}}$. Moreover, by Lemma 5.9, ExtConComp is $(0,\delta^{\prime})$-DP w.r.t. $f^{\mathrm{RR}}$. Therefore, by Theorem 4.9 and the fact that the composition of $\mathrm{RR}_{\epsilon,\delta}$ and $\mathrm{RR}_{0,\delta^{\prime}}$ is $(\epsilon,\delta+\delta^{\prime})$-DP, ExtConComp is $(\epsilon,\delta+\delta^{\prime})$-DP w.r.t. the verification function $f^{*}=f^{\gamma_{1},\gamma_{2}}$, finishing the proof. ∎
Remark 5.12 (Improved Basic Composition).
An immediate consequence of Lemma 5.11 is an improvement in the well-known basic composition theorem, stating that the composition of $k$ randomized response mechanisms $\mathrm{RR}_{\epsilon_{1},\delta_{1}},\dots,\mathrm{RR}_{\epsilon_{k},\delta_{k}}$ is $\left(\sum_{i=1}^{k}\epsilon_{i},\sum_{i=1}^{k}\delta_{i}\right)$-DP. By separately composing $\mathrm{RR}_{\epsilon_{1},0},\dots,\mathrm{RR}_{\epsilon_{k},0}$ using the original basic composition and composing $\mathrm{RR}_{0,\delta_{1}},\dots,\mathrm{RR}_{0,\delta_{k}}$ using Lemma 5.9, we can conclude that the composition of $\mathrm{RR}_{\epsilon_{1},\delta_{1}},\dots,\mathrm{RR}_{\epsilon_{k},\delta_{k}}$ is $\left(\sum_{i=1}^{k}\epsilon_{i},1-\prod_{i=1}^{k}(1-\delta_{i})\right)$-DP, where $1-\prod_{i=1}^{k}(1-\delta_{i})\leq\sum_{i=1}^{k}\delta_{i}$ always holds.

### v4

@@ 3. Continual Mechanisms
Let $\mathcal{M}$ be a CM and $\mathcal{I}$ be an identifier for $\mathcal{M}$. A verifier $\mathcal{V}$ with a verification function $f$, denoted by $\mathcal{V}[f]$, stands between an adversary $\mathcal{A}$ and the interactive mechanism $\mathcal{I}\!\circ^{*}\!\mathcal{M}$. Upon receiving a message from $\mathcal{A}$, $\mathcal{V}[f]$ applies $f$ to verify whether the sequence of pairs sent by the adversary up to that point satisfies a set of properties. Comparing the sequences formed by the first and second elements of the pairs, $f$ can be used to encode a neighboring property. If the sequences satisfy the property, $\mathcal{V}[f]$ forwards $\mathcal{A}$’s message to $\mathcal{I}\!\circ^{*}\!\mathcal{M}$; otherwise, it halts the communication. The formal definitions of identifiers, verification functions, and verifiers are given below.
@@
Since the set of all possible messages does not exist, we assume that all message spaces considered in this paper are subsets of some universal set $\mathcal{U}$. For instance, in (haney2023concurrent, ), the universal set is taken to be $\{0,1\}^{*}$. The domain of verification functions is $\mathcal{U}^{*}$.
@@
For a verification function $f$, a verifier $\mathcal{V}[f]$ is an IPM with the initial state space $S_{\mathcal{V}[f]}^{\text{init}}=\{()\}$, where $()$ represents an empty sequence. Since the initial state space of $\mathcal{V}[f]$ is a singleton, we denote an instance of $\mathcal{V}[f]$ with the initial state $()$ as $\mathcal{V}[f]$. The IPM $\mathcal{V}[f]$ stands between two IMs, referred to as the left and right mechanisms. The state of $\mathcal{V}[f]$ consists of the sequence of messages it has received from the left mechanism so far. Upon receiving a message $m$ from the left mechanism, $\mathcal{V}[f]$ appends $m$ to its state and applies the verification function $f$. If the sequence is $f$-valid, then $\mathcal{V}[f]$ forwards the message $m$ to the right mechanism; otherwise, it halts the communication. $\mathcal{V}[f]$ forwards any message from the right mechanism unchanged to the left mechanism.
@@
Let $S_{\mathcal{M}}^{\text{init}}=\{s^{\text{init}}\}$ be the initial state space of the continual mechanism $\mathcal{M}$ in Definition 3.7. By definition, the initial state spaces of $\mathcal{I}$ and $\mathcal{V}[f]$ are $\{0,1\}$ and $\{()\}$, respectively. Thus, the initial state space of $\mathcal{V}[f]\!\circ^{*}\!\mathcal{I}\!\circ^{*}\!\mathcal{M}$ is $\left\{((),0,s^{\text{init}}),((),1,s^{\text{init}})\right\}$. Define $\sim_{01}$ as the universal relation on this set, where every two elements are neighboring. Comparing Definition 3.7 and Definition 2.13 leads to the following corollary:
@@ 4. Concurrent Composition of Continual Mechanisms
Concurrent Composition of IMs. Vadhan and Wang (vadhan2021concurrent, ) extended the definition of the composition of NIMs to the concurrent composition of IMs, where an adversary interacts with $k$ interactive mechanisms $\mathcal{M}_{1},\dots,\mathcal{M}_{k}$ at the same time. In concurrent composition, the adversary generates a message and determines the receiver of this message (one of the $\mathcal{M}_{i}$s) adaptively based on all the previous messages exchanged between the adversary and the mechanisms $M_{1},\dots,M_{k}$. To formalize this, Vadhan and Wang (vadhan2021concurrent, ) introduced an interactive mechanism ${\textrm{ConComp}}[\mathcal{M}_{1},\dots,\mathcal{M}_{k}]$, which internally runs $\mathcal{M}_{1},\dots,\mathcal{M}_{k}$ and, upon receiving a message $(q,j)$, forwards the query $q$ to $\mathcal{M}_{j}$ and returns its response.
@@
Vadhan and Wang (vadhan2021concurrent, ) show that if each IM $\mathcal{M}_{i}$ is $\epsilon_{i}$-DP w.r.t $\sim_{i}$, and if the composition of NIMs $\mathrm{RR}_{\epsilon_{1},0},\dots,\mathrm{RR}_{\epsilon_{k},0}$ is $(\epsilon,\delta)$-DP, then ${\textrm{ConComp}}[\mathcal{M}_{1},\dots,\mathcal{M}_{k}]$ is also $(\epsilon,\delta)$-DP against an adaptive adversary. Lyu (lyu2022composition, ) and Vadhan and Zhang (vadhan2022concurrent, ) extend this result for $(\epsilon_{i},\delta_{i})$-DP IMs.
In the concurrent composition of IMs, queries asked to each mechanism are chosen adaptively, not only based on that mechanism’s previous responses but also using the responses of the other mechanisms up to that point. The results of (vadhan2022concurrent, ) and (lyu2022composition, ) show that this concurrency and additional adaptivity do not compromise privacy guarantees. As a result, all $(\epsilon,\delta)$-DP composition theorems for NIMs, such as the basic (dwork2006calibrating, ), advanced (dwork2010boosting, ), and optimal (KairouzOV15, ) composition theorems, can also be applied to the concurrent composition of IMs.
@@
To extend the proofs of (lyu2022composition, ; vadhan2022concurrent, ) on the privacy guarantee of the concurrent composition of IMs to the concurrent composition of CMs, we want to use the following intermediate lemma from their paper: For every $(\epsilon,\delta)$-DP interactive mechanism $\mathcal{M}$ and every pair of neighboring initial states $s$ and $s^{\prime}$ for $\mathcal{M}$, there exists an IPM $\mathcal{P}$ such that for each $b\in\{0,1\}$, the IMs $\mathcal{P}\!\circ^{*}\!\mathrm{RR}_{\epsilon,\delta}(b)$ and $\mathcal{M}(b)$ are equivalent. However, as we discuss in Section 4.2, there are hidden assumptions in Lyu’s proof of this lemma.
@@
We extend the definition of ConComp in (vadhan2021concurrent, ) and introduce ExtConComp that allows the creation of new mechanisms in addition to queries to existing ones. A creation query $(\mathcal{M},s,\epsilon,\delta,f)$ is a message that requests the creation of a continual mechanism $\mathcal{M}$ with the initial state $s$ and asserts that $\mathcal{M}$ is $(\epsilon,\delta)$-DP w.r.t. the verification function $f$. The mechanism ExtConComp accepts creation queries and messages of the form $(q,i)$, where $q$ is a query for the $i$-th created mechanism. The formal definition of ExtConComp is as follows:
@@
Remark 4.2.
By Russell’s paradox, a set of all continual mechanisms cannot exist. Therefore, to ensure that Definition 4.1 is mathematically correct and the query and answer sets of ExtConComp are well defined, we restrict creation queries to select mechanisms from a well-defined set and verification functions from a well-defined set. These sets may be infinite. The mechanism ExtConComp itself may belong to the mechanism set and, consequently, may be recursively created by itself.
@@
Next, we formalize these shared constraints by defining suitable message sequences. Later, when defining specific verification functions $f$ for different concurrent compositions, we set $f(m_{1},\dots,m_{k})=\bot$ whenever the sequence of messages $m_{1},\dots,m_{k}$ is not suitable. For convenience, in the rest of the paper, we represent a pair of identical creation queries $(\mathcal{M},s,\epsilon,\delta,f)$ by $(\mathcal{M},s,\epsilon,\delta,f)^{2}$.
@@ 4.2. $(\epsilon,\delta)$-DP IMs and $\mathrm{RR}_{\epsilon,\delta}$
Lyu (lyu2022composition, ) and Vadhan and Zhang (vadhan2022concurrent, ) show that for every $(\epsilon,\delta)$-DP IM $\mathcal{M}$ and every pair of neighboring initial states $s_{0}$ and $s_{1}$, there exists an IPM $\mathcal{T}$ (with a single initial state) such that for each $b\in\{0,1\}$, the mechanisms $\mathcal{T}\!\circ^{*}\!\mathrm{RR}_{\epsilon,\delta}(b)$ and $\mathcal{M}(s_{b})$ are equivalent. Although not explicitly stated, the construction of $\mathcal{T}$ proposed in (lyu2022composition, ) requires the answer sets of $\mathcal{M}(s_{0})$ and $\mathcal{M}(s_{1})$ to be finite. Vadhan and Zhang (vadhan2022concurrent, ) make an even stronger assumption that the query set must be finite. We instead assume that the answer distribution is discrete, which is formally defined as follows:
@@
An IM $\mathcal{M}$ has discrete answer distributions if $A_{\mathcal{M}}$ is countable, and for every initial state $s$, integer $t\in\mathbb{N}$, history $(q_{1},a_{1},\dots,q_{t-1},a_{t-1})\in(Q_{\mathcal{M}}\times A_{\mathcal{M}})^{t-1}$, and new query $q_{t}\in Q_{\mathcal{M}}$, the distribution of $\mathcal{M}(s)$’s answer in response to $q_{t}$ conditioned on having the history $(q_{1},\dots,a_{t-1})$ has a probability mass function (PMF).
@@
Lyu (lyu2022composition, ) additionally assumes that there exists a pre-fixed upper bound $c$ on the number of queries $\mathcal{M}$ answers. This assumption stems from their use of a backward-defined control function. By employing the limit control functions which we introduce in Section 10, one can modify their proof to remove this additional assumption. Moreover, using the formalization of Section 10 to define recursive functions, we can relax the requirement that the answer set is finite to allow countably infinite answer sets.
@@
Lemma 4.7 (slight extension of (lyu2022composition, ; vadhan2022concurrent, )).
@@ 4.3. Concurrent Composition of CMs $\mathcal{M}_{1},\dots,\mathcal{M}_{k}$
In the concurrent composition of IMs, queries to each mechanism are chosen adaptively, not only based on that mechanism’s previous responses but also using the responses of the other mechanisms up to that point. By definition, the choice of neighboring initial states for the interactive mechanisms is non-adaptive. However, in the concurrent composition of CMs, the choice of differences between query sequences, i.e., the neighboring query sequences is adaptive, based not only on the responses of the corresponding mechanism but also on the responses of other mechanisms. (Note that, unlike IMs, CMs inherently have a single fixed initial state; hence, no choice of initial state is involved.) Despite this additional adaptivity, in the following theorem, we show that the privacy results of (lyu2022composition, ; vadhan2021concurrent, ) for ${\textrm{ConComp}}[\mathcal{M}_{1},\dots,\mathcal{M}_{k}]$ also hold for $\mathcal{V}[f^{\alpha_{1},\dots,\alpha_{k}}]\!\circ^{*}\!\mathcal{I}\!\circ^{*}\!{\textrm{ExtConComp}}$.
@@ 4.4. Concurrent Composition of CMs with Fixed Parameters
We refer to the $f^{\epsilon_{1},\delta_{1},\dots,\epsilon_{k},\delta_{k}}$-concurrent composition of CMs as the concurrent composition of CMs with privacy parameters $\left((\epsilon_{i},\delta_{i})\right)_{i=1}^{k}$. A slight adaptation of the proof of Theorem 4.10 implies the following theorem.
@@ 5.2. Composition Theorem
In the proofs of the previous composition theorems (namely, Theorems 4.10 and 4.12), each created CM $\mathcal{M}_{j}^{\prime}$ with privacy parameters $(\epsilon_{j}^{\prime},\delta_{j}^{\prime})$ is simulated by the IPM $\mathcal{T}_{j}$ from Lemma 4.7. For each $b\in\{0,1\}$, the mechanism $\mathcal{T}_{j}$ interacts with $\mathrm{RR}_{\epsilon_{j}^{\prime},\delta_{j}^{\prime}}(b)$ once and simulates $\mathcal{V}[f_{j}^{\prime}]\!\circ^{*}\!\mathcal{I}(b)\!\circ^{*}\!\mathcal{M}_{j}^{\prime}$ identically. Next, we present a new construction for $\mathcal{T}_{j}$, where it interacts with $\mathrm{IRR}_{\epsilon_{j}^{\prime},\delta_{j}^{\prime}}$ instead of $\mathrm{RR}_{\epsilon_{j}^{\prime},\delta_{j}^{\prime}}$. This construction satisfies additional properties needed in the proof of Theorem 5.4. In Section 10, we prove the following lemma, which directly implies the desired IPM as stated in Corollary 5.7.
@@
Since the parameters $\delta_{j}^{\prime}$ are chosen adaptively, the composition of $\mathrm{RR}_{0,\delta_{1}^{\prime}},\mathrm{RR}_{0,\delta_{2}^{\prime}},\dots$ is an instance of the filter composition of NIMs (rogers2016privacy, ). We do not state this composition in terms of filters and define it using a verification function $f^{\delta^{\prime}}_{\mathrm{RR}}$ that only allows the creation of randomized response mechanisms $\mathrm{RR}_{0,\delta_{j}^{\prime}}$ and returns $\bot$ if $1-\prod(1-\delta_{j}^{\prime})>\delta^{\prime}$.
@@
The following lemma is proved in Section 11.  [^11] By setting $\epsilon_{1}=\epsilon_{2}=\dots=0$ in Theorem 2 of (whitehouse2023fully, ), it follows that for every $\epsilon>0$, $\delta^{\prime}>0$, and $0\leq\delta^{\prime\prime}\leq 1$, the composition of randomized response mechanisms $\mathrm{RR}_{0,\delta_{1}},\mathrm{RR}_{0,\delta_{2}},\dots$ with adaptively chosen privacy parameters $\delta_{1},\delta_{2},\dots$ satisfying $\sum_{i}\delta_{i}\leq\delta^{\prime\prime}$ guarantees $(\epsilon,\delta^{\prime}+\delta^{\prime\prime})$-DP. Taking $\epsilon$ and $\delta^{\prime}$ to zero, we can conclude that this composition is $(0,\delta^{\prime\prime})$-DP. To get tighter bounds, we relax the requirement $\sum_{i}\delta_{i}\leq\delta^{\prime\prime}$ to $1-\prod_{i}(1-\delta_{i})\leq\delta^{\prime\prime}$ and prove the composition is still $(0,\delta^{\prime\prime})$-DP.
@@
For every $\epsilon\geq 0$ and $0\leq\delta,\delta^{\prime}\leq 1$, if ${\textrm{Comp}}(\mathrm{RR}_{\epsilon_{1},0},\dots,\mathrm{RR}_{\epsilon_{k},0})$ is $(\epsilon,\delta)$-DP, then the $f^{\epsilon_{1},\dots,\epsilon_{k}}_{\infty,\delta^{\prime},RR}$-concurrent composition of CMs is $(\epsilon,1-(1-\delta)(1-\delta^{\prime}))$-DP.
@@
It remains to prove that the property (2) holds. The assumption that ${\textrm{Comp}}(\mathrm{RR}_{\epsilon_{1},0},\dots,\mathrm{RR}_{\epsilon_{k},0})$ is $(\epsilon,\delta)$-DP is equivalent to that ExtConComp is $(\epsilon,\delta)$-DP w.r.t. the verification function $f^{\beta_{1},\dots,\beta_{k}}$. Moreover, by Lemma 5.9, ExtConComp is $(0,\delta^{\prime})$-DP w.r.t. $f^{\mathrm{RR}}$. By the optimal composition theorem of (murtagh2015complexity, ), we have that the composition of $\mathrm{RR}_{\epsilon,\delta}$ and $\mathrm{RR}_{0,\delta^{\prime}}$ is $(\epsilon,1-(1-\delta)(1-\delta^{\prime}))$-DP. [^12] Replacing $k=2$, $\epsilon_{1}=\epsilon$, $\delta_{1}=\delta$, $\epsilon_{2}=0$, $\delta_{2}=\delta^{\prime}$, and $\epsilon_{g}=\epsilon$ in Theorem 1.5 of (murtagh2015complexity, ) implies that for every $\delta_{g}$ such that $0\leq 1-\frac{1-\delta_{g}}{(1-\delta)(1-\delta^{\prime})}$, the composition of $\mathrm{RR}_{\epsilon,\delta}$ and $\mathrm{RR}_{0,\delta^{\prime}}$ is $(\epsilon,\delta_{g})$-DP. $\delta_{g}=1-(1-\delta)(1-\delta^{\prime})$ satisfies this inequality. Therefore, by Theorem 4.10, ExtConComp is $(\epsilon,1-(1-\delta)(1-\delta^{\prime}))$-DP w.r.t. the verification function $f^{*}=f^{\gamma_{1},\gamma_{2}}$, finishing the proof. ∎